OpenStudy (anonymous):

Made up equation 1: 26-16y(19+4)=8y-32..... Is this equation able to be solved or not? If yes, give the answer and why. If not, tell why it isn't.

OpenStudy (lukecrayonz):

Yes it can be solved. y=29/188

OpenStudy (precal):

26-16y(23)=8y-32 26-368y=8y-32 26-376y=-32 -376y=-58 y=29/188